EG012B

PI: Gurvits
Title: High Z quasars

Session: Oct 1996 (session 3)
Frequency: 6 cm
Mode: B
Obs time (UT): 301 EU 13:30(27/10)-08:00(28/10)
Stations: EbJbNtToShUrWb
Processor: Bonn
EU Mk3 Tapes/St.: 4

Comment


To give feedback on this experiment, click here

Experiment Feedback

Torun (Roman Feiler at Mon Oct 28 19:20:28 EST 1996) Minor failures
2 minutes of the 4th pass lost due to low tape sensor failure

Effelsberg (D Graham at Tue Oct 29 02:43:37 EST 1996) Success
no known problems

HartRAO (Jonathan Quick at Tue Oct 29 04:59:06 EST 1996) Success
No known problems

Westerbork (tony foley at Tue Oct 29 09:10:55 EST 1996) Minor failures
Looks OK, but field system froze for the final minute, so the last tape had to be stopped by hand.

Noto (Corrado Trigilio at Wed Oct 30 06:02:42 EST 1996) Success
No problems reported. Cloudy sky.

Shanghai (Huang Xinyong at Tue Nov 5 03:48:38 EST 1996) Success
No problems, clock offset = 0.0 micro sec, and weather was cloudy.

Jodrell (Alastair Gunn at Thu Nov 7 07:14:05 EST 1996) Minor failures
Lost control of telescope for part of scan at 16:21 UT. Telescope sent to 1557+031 for next pass. Tapes sent to Bonn on 28/10/96 via Woodcroft. Weather wet and windy.

Urumqi (Liu Xiang at Tue Nov 19 03:41:15 EST 1996) Success
Clock offset 0.1+/-0.4us.

Simeiz (Alexandre Volvatch at Tue Nov 19 03:56:06 EST 1996) Minor failures
10 passes (14:24,16:39-17:17,04:04-05:25) were omited because we had not power (switch off).

MPIfR (Scott Aaron at Tue Sep 23 14:29:58 METDST 1997) Minor failures
Standard Westerbork problem. Rest OK.